# AnyBipe: An End-to-End Framework for Training and Deploying Bipedal Robots Guided by Large Language Models

### v1

@@ I Introduction
With the integration of advanced algorithms, enhanced physical simulations, and improved computational power, robotics has made significant strides[1, 2]. These innovations enable robots to perform tasks ranging from industrial automation to personal assistance with unprecedented efficiency and autonomy[3, 4]. As industrial robotics matures, researches focus on humanoid robots, particularly in replicating human-like characteristics and enabling robots to perform human-oriented tasks[5]. Correspondingly, bipedal robots can be used to simulate the morphology of human lower limbs, thus providing a method to explore the locomotion skills of humanoid robots [6, 7].
@@
Despite these advancements, training and deploying RL algorithms remains challenging. Effective reward design requires careful consideration of task-specific goals and the incorporation of safety constraints for real-world applications[10, 11]. This complexity demands significant engineering effort in training, testing, and iterative refinement. Although reward shaping and safe RL offer potential solutions[12, 13], they often rely on prior experience, complicating the reward design process. Furthermore, bridging the gap between simulations and real-world conditions—the ”Sim-to-Real” challenge—remains difficult[14]. Techniques such as domain randomization[15], which randomizes physical parameters to enhance agent robustness, and observation design, which facilitates task transfers across varied terrains, prove formidable, but still require real-world testing and human feedback for model selection and fine-tuning[7, 16, 17].
Hohe integration of large language models (LLMs) into robotics offers the prospect of reducing human work in such process. Known for their capabilities in code generation and missing planning, LLMs are being widely applied to complex robotics applications[18]. For instance, they play a pivotal role in embodied intelligence by enabling the dynamic creation of action tasks[19]. Recent developments have explored the utility of LLMs in improving reward function design and refining from simulation performance—key areas that reduce the need for human intervention[20]. However, a comprehensive framework that designs suitable policies and automatically implements properly trained models in real-world remains lacking. To address this issue, we propose a novel framework that leverages LLMs to optimize the entire training-to-deployment process. The framework minimizes human involvement, enables autonomous design, training, and deployment of RL algorithms, and supports exploring strategies from scratch or enhancing existing ones.
@@
This paper is organized as follows: First, we introduce the design principles and implementation methods of the framework modules in detail. Then, we describe the experiments used to verify the role of each module and express the training and deployment effects of the framework. Finally, we summarize our contributions and propose future research directions for our framework.
@@ II Related Works
Large Language Model Guided Robotics. Large language models (LLMs) have demonstrated considerable capabilities in task understanding[36, 37], semantic planning[38, 39], and code generation[40, 41, 42], allowing them to be effectively integrated into a variety of robotic tasks. LLMs automate environmental analysis[43, 44], reward functions design[45, 46], and task-action mappint[47, 48]. However, challenges such as data scarcity, real-time performance, and real-world integration remain[49]. In addition, due to the lack of awareness of actual data, LLM-driven code design sometimes needs to be optimized through human feedback. Our framework addresses the existing limitations by adopting environmental characteristics and safety constraints as priors, and uniquely combines homomorphic feedback from real-world applications, reducing the need for human feedback in the process.
Sim-to-real Training and Deploying Techniques. The gap between simulated environments and real-world conditions, known as the “reality gap,” presents significant challenges for deploying RL strategies in robotics[50, 51]. Techniques such as domain randomization[52, 53, 54] and system identification[55, 56, 57] are widely used to address this issue. Researchers have proposed sim-to-real solutions for bipedal robots to handle tasks such as turning and walking[58, 59]. Recent work has also integrated LLMs to enhance environmental modeling and reward function design, making simulations more reflective of real-world complexity[20]. However, most approaches still rely on separate training in simulation and real-world evaluation, often using human feedback to assess sim-to-real effectiveness. Our work extends these techniques by introducing an evaluation loop that continuously monitors sim-to-real performance during deployment.
@@ III Methods
In this section, we introduce the AnyBipe framework in detail. The framework consists of three modules aiming to automate reward design, training and simulation, and deployment feedback to minimize the need for manual work. The framework requires the robot’s URDF model, a basic task description, an RL training platform, and a ROS-based system for communication and control [60], along with an SDK for sensor and actuator data, as compulsory input. Optional elements include a manual reward function template, a teacher model realized with previous work, and custom environmental observations for evaluation. AnyBipe generates reward functions accordingly and trains policies supervised by the teacher model. A group of top strategies, determined by evaluation criteria and simulation results, are deployed via ROS. After validating in the ROS Gazebo simulation, policies that pass the safety check undergo real-world testing, and the best strategy is then selected. Evaluation result across various environments and the selected best policy are compiled to feedback, guiding iterative improvements in reward generation. The procedural steps are encapsulated in Algorithm 1.
III-A Module 1: LLM Guided Reward Function Design
@@
To address key issues, we developed a robust context establishment mechanism that effectively tackles underdesigned reward functions and safety constraints. Our approach classifies prompts into two categories: General and Task-Specific. For general tasks, we provide coding tips, function templates, and predefined templates that facilitate code compilation, training feedback, and evalu feedback. We also integrate reference tables from Isaac Gym for environmental variables like motor torque, torque limits, and feet height. These tables prevent LLMs from the customization of non-existent observations while enhancing the utilization of environmental variables in in order to ensure that LLMs consider actionable constraints during reward function design. Our experimental results affirm that LLMs can seamlessly incorporate these safety restrictions and environmental variables, thus designing highly effective reward functions, evidencing their exceptional context tracking and directive following capabilities [62, 63].
Furthermore, the Task-Specific module allows users to define custom prompts for specific tasks, facilitating the rapid generation of viable code and standardization of reward calculations. Users have the flexibility to use trainable artificial reward functions as templates or employ various computational paradigms to enhance the accuracy and applicability of reward assessments.
To refine the evaluation of designed reward functions for completing the improvement loop, we introduced a comprehensive reward function evaluation scheme. This scheme not only tracks changes in rewards and observations throughout the training but also integrates a homomorphic evaluation model to closely assess real-world robot performance. This model ensures a high correlation between real-world outcomes and theoretical reward functions, enabling LLMs to intuitively identify the most impactful components of the reward functions. Details on this model are expounded in subsection III-C, showcasing our commitment to precise and practical feedback mechanisms that enhance real-world applicability.s more significantly to real-world performance.
III-B Module 2: RL Training Adopting Reference Policy as Teacher
In this section, we detail the adaptations applied to guide RL training towards desired actions using frameworks provided by Legged Gym222https://github.com/leggedrobotics/legged_gym, employing the Proximal Policy Optimization (PPO) algorithm[64] as our foundation. We assume the existence of a baseline policy , which could be derived from traditional control methods or previous RL techniques.
@@ III-B Module 2: RL Training Adopting Reference Policy as Teacher
The approximation uses the integral properties of the Dirac function and expression of KL divergence, which is omitted due to space constraints. With adequate observations, this approximation becomes reliable. This improvement can effectively prevent reinforcement learning from falling into a degradation trap or a local optimal solution.
In the AnyBipe framework, we have established a template for deploying existing policies as teacher functions, demonstrating integration of models deployed using PyTorch, ONNX, and traditional control algorithms implemented in C++ into our framework. This will allow users to transfer previous work to current work, or introduce pre-trained policies for simple tasks to accelerate the convergence of policies on complex tasks.
III-C Module 3: Deployment, Evaluation, and Feedback
In this section, we discuss optimal strategy selection, state estimator implementations, and methods to align evaluation with reward functions, leading to refined strategy generation by LLMs.
To shield our data from ineffective reward functions, we define essential training metrics—like command compliance and survival duration—as basic measurements , filtering samples that meet these benchmarks. For specific tasks such as bipedal stair climbing or standing, customization of observational metrics and coefficients is enabled in configuration files. We then select the top 15% of strategies for real-world deployment, prioritizing those that maximize combined metrics:
Strategies are deployed in a ROS-based robotic framework () using ONNX. Given that the training () and real environments () are isomorphic, we define a homomorphism , ensuring the real-world evaluation metric mirrors the reward function. Our automated script aligns reward structures with observed real-world data, informing users of any mismatches.
We also develop a basic state estimator for comprehensive robotic assessments, including metrics like step width and leg height. Models are first validated in Gazebo to confirm operational stability, with safety checks aligned with safety regulations proposed in subsection III-A. Successful models undergo real-world testing, with optimal models selected based on:
AnyBipe enables an autonomous cycle from reward function generation to training, deployment, and optimization, facilitating user-driven training of bipedal robot RL algorithms without pre-existing strategies.
@@ IV Experiments
Our experiments were conducted on a six-degree-of-freedom bipedal robot from Limx Dynamics, shown in Fig. 4. GPT-4o was selected as the LLM base, with training performed on an NVIDIA GTX 3090Ti. We explored the robot’s locomotion on both flat and complex terrains over multiple experimental rounds. This section highlights key experiments on individual modules and overall training effectiveness. Due to space limitations, only critical results are presented here, with detailed data and charts available on our GitHub page.
Table I outlines key reward functions that maintain the same structure across both human-engineered and AnyBipe-generated rewards, differing only in scale, allowing direct performance comparison.
@@
IV-A Module Analysis
We evaluated each module by testing prompt design and various augmentations using gpt-3.5-turbo, gpt-4, and gpt-4o with samples per batch. Success rates—defined as the proportion of samples processed successfully in the first iteration—were compared across four configurations: basic prompts, prompts with code references, safety regulations, and both. Fig. 5 shows that adding code references and safety regulations improves LLM performance in generating usable reward functions. As gpt-4 and gpt-4o show similar performance with complete prompts, gpt-4o is recommended as the primary model.
We validated the effect of incorporating safety regulations into reward functions by comparing models with and without safety prompts in Isaac Gym and Gazebo environments. Fig. 6 shows the operational postures and IMU spatial angles of safe models (c)(d) versus unsafe models (a)(b). Results indicate that, even without real-world feedback, safety prompts effectively constrain robot behavior, supporting practical deployment.
We evaluated teacher-guided models under artificial reward functions, training each for 5,000 iterations on complex terrain. The original model used only an artificial reward function, while the teacher-guided model was instructed by an operational ONNX model with . Performance was measured using reward success and terrain level. Results showed that the teacher-guided model had more stable training, with faster and less volatile reward growth.
For the Homomorphic evaluation part, the following table shows some evaluation indicators before and after the conversion, and the evaluation results under the best model. It can objectively reflect some differences from simulation to reality. Each column seperately represents reward function name, reward in isaac gym, homomorphic measurement in gazebo, in reality, and the mapped tracking result in reality (30 seconds of tracking).

### v2

@@ I Introduction
With the integration of advanced control algorithms, enhanced physical simulations, and improved computational power, robotics has made significant strides[1, 2]. These innovations enable robots to perform tasks ranging from industrial automation to personal assistance with unprecedented efficiency and autonomy[3, 4]. As industrial robotics matures, researches focus on humanoid robots, particularly in replicating human-like characteristics and enabling robots to perform human-oriented tasks[5]. Correspondingly, bipedal robots can be used to simulate the morphology of human lower limbs, thus providing a method to explore the locomotion skills of humanoid robots [6, 7].
@@
Despite these advancements, training and deploying RL algorithms remains challenging. Effective reward design requires careful consideration of task-specific goals and the incorporation of safety constraints for real-world applications[10, 11]. This complexity demands significant engineering effort in training, testing, and iterative refinement. Although reward shaping and safe RL offer potential solutions[12, 13], they often rely on prior experience, complicating the reward design process. Furthermore, bridging the gap between simulations and real-world conditions—the “Sim-to-Real” challenge—remains difficult[14]. Techniques such as domain randomization[15], which randomizes physical parameters to enhance agent robustness, and observation design, which facilitates task transfers across varied terrains, prove formidable, but still require real-world testing and human feedback for model selection and fine-tuning[7, 16, 17].
The integration of large language models (LLMs) into robotics offers the prospect of reducing human work in such process. Known for their capabilities in code generation and missing planning, LLMs are being widely applied to complex robotics applications[18]. For instance, they play a pivotal role in embodied intelligence by enabling the dynamic creation of action tasks[19]. Recent developments have explored the utility of LLMs in improving reward function design and refining from simulation performance—key areas that reduce the need for human intervention[20]. However, a comprehensive framework that designs suitable policies and automatically implements properly trained models in real-world remains lacking. To address this issue, we propose a novel framework that leverages LLMs to optimize the entire training-to-deployment process. The framework minimizes human involvement, enables autonomous design, training, and deployment of RL algorithms, and supports exploring strategies from scratch or enhancing existing ones.
@@
This paper is organized as follows: First, we introduce the design principles and implementation methods of the framework modules in detail. We then describe the experiments used to verify the role of each module and express the training and deployment effects of the framework. Finally, we concluded our contributions and propose future research directions for our framework.
@@ II Related Works
Large Language Model Guided Robotics. Large language models (LLMs) have demonstrated considerable capabilities in task understanding[36, 37], semantic planning[38, 39], and code generation[40, 41, 42], allowing them to be effectively integrated into a variety of robotic tasks. LLMs automate environmental analysis[43, 44], reward functions design[45, 46], and task-action mapping[47, 48]. However, challenges such as data scarcity, real-time performance, and real-world integration remain[49]. In addition, due to the lack of awareness of actual data, LLM-driven code design sometimes needs to be optimized through human feedback. Our framework addresses the existing limitations by adopting environmental characteristics and safety constraints as priors, and uniquely combines homomorphic feedback from real-world applications, reducing the need for human feedback in the process.
Sim-to-real Training and Deploying Techniques. The gap between simulated environments and real-world conditions, known as the “reality gap”, presents significant challenges for deploying RL strategies in robotics[50, 51]. Techniques such as domain randomization[52, 53, 54] and system identification[55, 56, 57] are widely used to address this issue. Researchers have proposed sim-to-real solutions for bipedal robots to handle tasks such as turning and walking[58, 59]. Recent work has also integrated LLMs to enhance environmental modeling and reward function design, making simulations more reflective of real-world complexity[20]. However, most approaches still rely on separate training in simulation and real-world evaluation, often using human feedback to assess sim-to-real effectiveness. Our work extends these techniques by introducing an evaluation loop that continuously monitors sim-to-real performance during deployment.
@@ III Methods
In this section, we introduce the AnyBipe framework in detail. The framework consists of three modules that collectively aim to enhance reward design, facilitate semi-supervised RL training, and automate evaluation and feedback. Users are required to prepare the URDF model, task description, basic RL environment, and an operational control platform such as ROS[60]. A runtime data collector, known as the State Tracker, which gathers essential runtime data such as IMU data, joint positions, velocities, and torques, is also essential. Starting with a well-engineered prompt framework for code generation, we enable LLMs to write appropriate reward functions based on user-provided patterns, while considering their functionality and potential safety issues. The generated reward functions are then trained in modified RL environments, aided by cold start techniques to improve convergence. The top-performing policies are subsequently validated using a homomorphic estimator to assess their safety and compare their performance with the training stage. All estimation results are then compiled into feedback prompts for the next iteration’s improvement. The procedural steps are encapsulated in Algorithm 1.
III-A Module 1: LLM Guided Reward Function Design with Proper Prompt Engineering
@@
To address these challenges, we have developed a robust context establishment mechanism to assist LLMs in constructing reliable and secure reward functions. As outlined in the algorithm, we employ an iterative ReAct (Reason + Act) framework [62] to improve the coding results based on real-world performance. The prompts are structured into three distinct phases: the system prompt stage, the task specification (initial user) stage, and the feedback (user) stage. The system prompt stage is dedicated to generating knowledge prompts [63] for the RL environment, as well as utilizing few-shot prompting [64] for reward function generation. Relevant system variables, such as “root states,” “feet height,” and “rigid body state,” are extracted from the environment and cataloged into a reference table to prevent the LLM from inventing non-existent variables. This approach also improves the integration of environmental variables within the set . For more intricate tasks, such as humanoid walking with a desired pattern, we encourage users to incorporate custom prompts following our “coding references” guide. This enables LLMs to learn from human-engineered reward functions and autonomously generate task-specific code. Additionally, a section on “coding restrictions” is provided as negative prompts, urging LLMs to account for safety constraints such as contact forces, degrees of freedom (DoF) limits, and torque limitations. Our experimental findings confirm that LLMs can effectively incorporate these safety restrictions and environmental variables, thereby designing highly efficient reward functions. These results underscore the models’ outstanding ability to track contextual information and adhere to directives [65, 66].
The initial task-specific stage serves to provide essential information for training, including a description of the robot, key parameters such as base height and average foot air time, as well as an introduction to reward-writing techniques. Following the first iteration, this stage is replaced by a comprehensive reward function evaluation scheme, detailed in subsection III-C, which functions as a feedback prompt. This prompt incorporates multiple components, including the performance of the reward function, the consistency of the homomorphic estimator with the reward functions, and whether the policy adheres to safety restrictions. The evaluator integrates pre-defined human preferences, such as human factor observation () and user-defined observation terms labeled with “gt” to maintain consistency across LLM-generated reward functions. The success rate is also considered to identify the best-performing sample, which is then fed back to guide the LLM in adjusting terms and coefficients.
Additionally, we incorporate self-consistency techniques [67] to assist the LLM in assessing the performance of individual reward terms, determining whether they exhibit good or poor performance, and making necessary modifications. In cases where the code fails to execute, error logs related to the reward file are extracted and provided as feedback. This stage also offers adjustment tips to further refine the reward function. We also guide the LLMs to generate reward functions in the form of Chain-of-Thought (CoT) [68], providing a step-by-step explanation of their implementation. The generated code is instructed to be presented within a code block at the end of the output, which is subsequently processed by our script to remove the tags and extract the target code within the code block. Our experimental results demonstrate that utilizing the CoT approach enhances the LLM’s ability to accurately identify and address issues within the code implementation.
III-B Module 2: RL Training Adopting Reference Policy for Cold-Start
In the initial phase of training, where Large Language Models (LLMs) generate rewards without external feedback, the reward terms often exhibit poor scaling, which can result in slow convergence—a challenge commonly referred to as the “cold start” problem. To address this issue, the integration of a reference policy can provide guidance, ensuring that the policy network converges toward the desired objective. This section outlines the adaptations made to direct reinforcement learning (RL) training toward the target actions, utilizing frameworks from Legged Gym and leveraging the Proximal Policy Optimization (PPO) algorithm [69] as the foundational approach. We assume the existence of a baseline policy, denoted as , which is derived from pre-existing functional policies. Interfaces are provided to accommodate three types of teacher inputs: policies implemented using traditional control techniques such as Model Predictive Control (MPC) or Whole Body Control (WBC), previously trained policies in formats like PyTorch or ONNX, and simple locomotion distributions, such as sine waves.
@@ III-B Module 2: RL Training Adopting Reference Policy for Cold-Start
With adequate observations, this approximation becomes reliable. The weight of this term should be set to a relatively small value in order to guide the policy net out of local minimum, but not preventing it from exploring better solution.
In the AnyBipe framework, we have developed a template for deploying existing policies as teacher functions, illustrating how such policies can be integrated into the reinforcement learning (RL) training process. The integration mechanism operates as follows: if a human-engineered reward function is designated as the ground-truth reward, it is trained prior to the iteration, and the trained policy is subsequently set as the teacher. Alternatively, users can provide their own implemented policies following our template. After each iteration, the reference policy is updated to the trained policy with the best performance if it passes a safety check. This metric compares the performance of human-engineered policies with those generated by Large Language Models (LLMs), offering essential insights into the correctness of metric implementation. Furthermore, it safeguards against the degradation of the LLM-generated reward function by tracking how the behavior of reward function terms evolves after LLM modification, following the same Markov Decision Process (MDP) path.
III-C Module 3: Feedback From Simulation and Deployment Stage With Minimal Human Effort
Traditionally, Eureka-like algorithms only collect training feedback for the subsequent generation, with generated policies left to engineers for examination and deployment. In contrast, our framework automates the entire process using Python, Bash, and C++ scripts, establishing a robust safety check criterion and providing a numerical measure for the sim-to-real gap, referred to as the “homomorphic estimator”. This estimator enables the Large Language Model (LLM) to identify which terms have a significant impact on deployment procedures.
To establish a general criterion for evaluating policy performance, we define a set of observations with “ground-truth” (gt) labels, which remain consistent across all generations. For all our experiments, this set includes linear velocity tracking results, angular velocity tracking results, and survival time. The product of these three indicators is considered the base factor, termed the environment estimator , which gauges the success of policy execution. Additionally, we provide users the flexibility to define custom observation/reward factors, termed human factors , within the configuration file, along with their corresponding weights . The weighted sum of these factors is then used to compute the total score, with policies ranked accordingly to select the best candidates for simulation and deployment validation.
We automatically deploy the policy file, in either JIT or ONNX format, to a user-specified simulation environment such as Gazebo or MuJoCo. The simulation script runs for at least , as defined in the OnPolicyRunner under the rsl_rl library, ensuring that the simulation duration matches or exceeds a full training epoch. Data from the robot, including IMU readings, fall-off time (or maximum if no fall-off occurs), joint speed, velocity, and torques, are collected for evaluation by our homomorphic evaluator. Given that the training () and real-world environments () are isomorphic, we define a homomorphism , ensuring that the real-world evaluation metric mirrors the reward function. Users are required to provide a mapping table linking IsaacGym system variables to simulation-tracked state variables (e.g., root_state in IsaacGym to part of qpos in MuJoCo). Fig .3 illustrates this conversion procedure, where the reward functions are transformed into new estimators that maintain consistency between training and simulation. The evaluator then provides a quantitative measure of the sim-to-real gap, identifying specific reward terms that may require improvement, thus assisting the LLM in refining its approach.
To ensure the policy is safe for real-world deployment, selected policies must pass a safety check, denoted as , before being compiled by ROS. Users are required to implement following the template provided. evaluates simulation data, including torques, positions, ground contact forces, and other relevant factors. If all checks pass, returns true, enabling the script to automatically compile and execute the reality deployment code (this stage can be passed if no reality robots are available). If any checks fail, the script identifies the failed terms and generates a feedback prompt. The reality evaluator operates similarly to the simulation evaluator, gathering and evaluating data. Users are allowed to monitor the process and can stop the deployment if dangerous behavior is detected. External stops are recorded by the script, indicating that the reality deployment has failed; otherwise, the deployment is considered successful. These steps also provide a score that contributes to policy selection, with the best generation calculated as follows:
Except for the final reality deployment stage, which is recommended to have a supervisor (though it can still operate without human oversight, or with VLM and cameras substituting human functions), the entire cycle operates autonomously. Evaluation, safety check, and deployment procedure are working out human-free. Even the initial reference rewards or pre-existing policies are optional, as LLMs demonstrate significant potential in handling zero-shot generation tasks [70]. This framework offers a method for automatically adjusting or tuning reward functions for robot engineers, while also providing a solution for developing RL policies from scratch for robots that have no prior RL implementations.
@@ IV Experiments
IV-A Experimental Setup
Our experiments were conducted with three types of bipedal robots: the 6 DoF pointfoot robot P1 from Limx Dynamics, the 10 DoF humanoid Unitree H1 (lower body only), and the 29 DoF humanoid from Turin Robots (12 lower body DoF, with other DoF fixed during training), as shown in Fig. 4. Each robot was equipped with different initial prompt sets: P1 used human-engineered reward functions as the ground truth for comparison; Unitree H1 used some human-engineered functions as templates; and Turin, with no prior RL implementation, utilized templates partially adapted from the open-source humanoid-gym project, with no ground-truth reward or reference policy. LLMs used include OpenAI GPT-4o[71], Anthropic Claude-3.5-Sonnet [72], and Deepseek-R1 [73], to validate the versatility of our prompt groups and scripts. The framework was trained on a GPU server with 4 NVIDIA RTX 3090 GPUs. The robots were trained on common and trimesh terrains, with task settings including iterations, samples, and the top (3 best samples) selected for each task. Thanks to the implemented GPU allocator algorithm, we are able to complete a run with 80 trainings in total in 8 hours, including LLM generation and evaluation time. For 2 humanoids, the framework functions with only simulation evaluation step with feedback. For P1, we are able to perform a real-world deployment test to validate our auto-deployment process. The GPU server is hardwared to connect to robot controllers, and simulation environments (MuJoCo mjviewer and Gazebo) are set to function under headless mode.
Table I outlines general environment estimators in the form of reward functions. These estimators maintains the same form across all training and evaluating process.
@@ IV-A Experimental Setup
IV-B Module Analysis
Before introducing the outcome for the whole framework process, we first conduct several experiments seperately on each module to examine its effectiveness. We evaluated the LLM performance, RL-training performance, homomorphic performance, and whether criterion functions normally.
We evaluated each LLM base by testing prompt design and various augmentation. The temperature for generation is set to 0.4 for all models. We define success rate (number of code that can be executed for all generation, SR) and max normalized reward success (max , the average on three experiments, Max S.) as two important indicators. The test result among the LLMs are shown as follows. We noticed that Deepseek-R1 and Claude-3.5-Sonnet shows best performance.
Name | P1 | H1 | Turin | SR | Max S. |
GPT-4o | 228/240 | 173/240 | 52/240 | 62.9% | 0.824 |
Claude-3.5 | 232/240 | 213/240 | 201/240 | 89.7% | 1.032 |
Deepseek-R1 | -/- | 232/240 | 211/240 | 92.3% | 1.060 |
To show the strength of reference policy, we take the training of P1 as example, set the reference policy using human-engineered reward function trained policy (5000 epochs training, flat terrain), and compare the training result using same reward function under trimesh terrain. The teacher coefficient is set to with curriculum learning on. Performance was measured using reward success and terrain level. Results in Fig. 6 showed that the teacher-guided model had more stable training, with faster and less volatile reward growth.
For the Homomorphic evaluation part, the following table shows some evaluation indicators before and after the conversion for P1 training, and the evaluation results under the best model. It can objectively reflect some differences from simulation to reality. Each column seperately represents reward function name, reward in isaac gym, homomorphic measurement in gazebo, in reality, and the mapped tracking result in reality (30 seconds of tracking, where env configuration requires at least 1.6s tracking).